obtain the differential equation of Asinx^2+Bcosx^2

Grade:12th pass

1 Answers

Debolina Sarkar
6 years ago
d/dx(Asinx^2+Bcosx^2)=A(d/dx sinx^2)+ B(d/dx cosx^2)[where A and B are two constants]We can clearly see that this problem each contains the composition of two functions Like, we can express Asinx^2=f(g(x)) Where g(x)=x^2; f(x)=sinxSimilarly Bcosx^2In this sort of problem we has to differentiate​ both functions and multiple together. This is Known as chain rule.By using the chain rule=Acosx^2*2x-Bsinx^2*2x